Consider the vector field equation Consider the vector field equation . Sketch t .

Sketch this vector field using arrows in the x-y coordinates in the range of -4<=x<=4 and -4<=y<=4.

Compute the line integral Consider the vector field equation . Sketch t from (x,y) = (0,1) to (1,) along the path of Consider the vector field equation . Sketch t . (Hint: Come up with a single parameter, such as t or Consider the vector field equation . Sketch t that describes the x and y components of the path.
Answer #1

F.dr=-x-y=-1+y2-y

integral of F.dr=-y+y3/3-y2/2|10

F=1-1/3-1/2=1-5/6=1/6